OpenStudy (anonymous):

A batter hits a baseball upward with an initial speed of 96 feet per second. After how many seconds does the ball hit the ground? Use the formula h = rt - 16t2 where h represents height in feet and r represents the rate in feet per second.

OpenStudy (mertsj):

\[0=96t-16t^2\]

OpenStudy (mertsj):

\[-16t(t-6)=0\]

OpenStudy (mertsj):

\[t=6\]

OpenStudy (anonymous):

0 = 96ft/s * t - 16 t^2 16t^2 = 96t 16t = 96 t = 96/16 = 6

OpenStudy (mertsj):

The ball hits the ground after 6 seconds.

OpenStudy (anonymous):

At the time the ball hits the ground means h = 0 -> h = - 16t² + 96t = 0 -> 16 t ( -t + 6 ) = 0 => t = 6 sec for the ball to land ( omit t = 0 as start time)